OpenStudy (anonymous):

Use the equation y = 20 – 4.50x to predict how much money John will have left over if he rents 3 movies. In the equation x represents the number of movies and y represents the amount of money. Answer $16.50 $13.50 -$6.50 $6.50

OpenStudy (anonymous):

I am working on it.

OpenStudy (anonymous):

y = amount of money x = 3 = # of movies y=20-4.5x y=20-4.5(3) y=6.5 movies

OpenStudy (anonymous):

oops.. $6.5